# A Survey on Class-Agnostic Counting: Advancements from Reference-Based to Open-World Text-Guided Approaches

### v3

@@ III-B Reference-based Class-agnostic Counting
The seminal work about reference-based object counting is represented by [37]. Here, Ranjan et al. formulated object counting as a few-shot regression task, defining the primary goal, the first benchmark, and the pioneering methodology. Specifically, as already mentioned in previous sections and illustrated in Fig. 4, they defined the task objective as inferring the number of objects belonging to novel classes unseen during training, given an image and a few exemplar objects. Since the task was novel, the authors introduced the first dataset tailored for it, a collection of images containing objects belonging to 147 categories named FSC-147; labels comprise dots for the training stage, and exemplar bounding boxes exploited both in the supervised learning and at inference time. See also Sec. IV for more details. Finally, the authors proposed the Few Shot Adaptation and Matching Network (FamNet), a novel architecture designed to address this task. Specifically, it includes two main modules: (i) a multi-scale feature extraction module and (ii) a density prediction module. The first consists of the first four blocks picked up from an ImageNet pre-trained ResNet-50 backbone [87], frozen during training. The convolutional feature maps produced by the third and fourth blocks provide the representation of the images $I_{i}$; furthermore, they are also exploited to obtain the multi-scale features of the exemplar images obtained from $B^{E}$ by performing ROI pooling [88] at different scales. Rather, the density prediction module consists of some stacked convolution blocks and upsampling layers; the last one is a $1\times 1$ convolutional layer in charge of predicting the final $2D$ density map $D^{map}$ having the same size as the input image. The input of this second module does not correspond to the features produced by the feature extractor, as usual in the class-specific counting task. Instead, the authors computed the convolution operation between the scaled exemplar features and the image features, obtaining multiple correlation maps, one for each scale, that are then concatenated and fed to the density predictor. The reason behind this is making the density regressor agnostic to the visual categories. During the learning stage, FamNet is trained by minimizing the mean squared error (MSE) between the predicted and the ground truth density map. Finally, this work also proposed a sort of fine-tuning that exploits test data to improve performance further. Specifically, it introduced an adaptation loss that exploits the locations of the exemplars included in the test set to be used only at inference time. It comprises two loss components: (i) the Min-Count loss derived from the intuition that the pixel sum of a predicted density map region corresponding to a bounding box defining an exemplar must be $\geq 1$, and (ii) the perturbation loss that essentially ground to the idea that density values around the location of exemplars in the predicted density map should ideally look like a Gaussian, similar to correlation filters in tracking algorithms [89, 90, 91].
@@
Roughly simultaneously, Yang et al. introduced a similar solution [58] named Class-agnostic Few-shot Object Counting Network (CFOCNet). The architecture follows the same scheme as FamNet, comprising an encoder and a decoder. The former has two streams for feature extraction, one for the query image and one for the exemplars (here referred to as reference images). It is based on the first three blocks of an ImageNet pre-trained ResNet-50 model [87]. The features concerning the exemplars are matched by computing the convolution operation with the output of a self-attention mechanism [92] applied to the features relating to the query image. These resulting matching maps are finally fused by exploiting a scale-aware fusing mechanism, and the output is fed into the decoder to compute the density map. The decoder is similar to the one exploited in FamNet – some stacked convolution blocks and upsampling layers – and it is supervised by standard MSE loss as well as by SSIM loss [93] to catch local pattern consistency [94].
@@
Some of the same authors of [37] proposed an extension of FamNet in [59]. Here, they addressed the limitations due to the narrow size of the dataset on which the supervised learning of FamNet was bounded, and they proposed the Vicinal Counting Network (VCN) [59]. The core of VCN is the same as FamNet with some slight changes; additionally, during training, the authors exploited a generator aiming at augmenting and synthesizing images from the vicinity of training samples, increasing the training set. Specifically, the generator is fed with an input image $I\in\mathbb{R}^{H\times W\times 3}$ and a noise vector $z\in\mathbb{R}^{m}$, and it produces an augmented version $I_{z}\in\mathbb{R}^{H\times W\times 3}$ of $I$. This augmented image serves as the input to the feature extractor and the subsequent regressor module, together with the original input $I$. The entire system is then trained jointly in a cooperative fashion. However, unlike the min-max game typical of GANs [95], VCN is trained by employing a standard minimization problem. Alongside the MSE loss for the density regressor optimization, the authors proposed a reconstruction loss in charge of forcing the generator to produce samples close to the original input samples and a diversity loss to avoid the generator learning identity mapping from $I$ to $I_{z}$. Finally, the authors used the same test time adaptation loss of FamNet [37].
@@
Another work inspired by FamNet is [60]. First, the authors explored the main limitations of FamNet by analyzing failure cases, and they found that counting performance is highly conditioned by the diversity of the considered exemplars, i.e., intra-class diversity. Specifically, exemplar boxes are limited in number – they potentially do not cover all object instances – and their diversity is not guaranteed since they are provided by annotators subjectively. To tackle this issue, Gong et al. introduced the Robust Class-Agnostic Counter (RCAC), designed to be more robust to intra-class diversity. The architecture follows a two-stream scheme. The first branch consists of a feature extractor similar to the one exploited in FamNet, which takes as input RGB images $I_{i}$. Conversely, the second branch consists of a different feature extractor based on the VGG backbone [96], which takes as input edge images obtained by processing $I_{i}$ throughout an edge detector. This second feature extractor is randomly initialized and, different from the RGB feature extractor, optimized during the training. The reason claimed by the authors behind using edge images is that edges are a sort of class-agnostic knowledge, i.e., shape is a more reliable cue across instances than color. Then, like in [37], ROI pooling is exploited for cropping exemplar box features from the feature maps generated by the two branches. Still, in this case, exemplar features are augmented with a novel feature augmentation module. After that, feature correlation is obtained by computing the convolution operation between the latter and the full feature maps, like in previous works, and, finally, the predicted density map is obtained with a density prediction module identical to the one introduced in FamNet. In detail, the feature augmentation module aims to generate more exemplars of different colors, shapes, and scales in the feature space by combining the provided exemplars. From this perspective, this module shares the same goal of the VCN [59], i.e., increasing the size of the training set.
@@
Another notable work that represents a milestone is [61]. Here, Shi et al. argued that two factors played crucial roles in previous works: (i) feature representation of query images and exemplars, and (ii) similarity metric for matching exemplar and image features. Previous methods either exploited a fixed [37, 59] or a learnable [58] feature extractor but employed a similarity metric relying on the fixed inner product [37, 59, 58, 60, 63], which yielded insufficient matching outcomes. Hence, the authors proposed the Bilinear Matching Network (BMNet+) to improve the similarity modeling. Specifically, the authors designed a learnable bilinear similarity loss to supervise the similarity-matching results, thus going beyond the fixed inner product. The authors took inspiration from metric learning, which aims to embed data into a space where similar samples are pulled close while dissimilar ones are pushed away [97]. In more detail, their similarity loss pulls close the features between the exemplar and target instances while pushing away the features between exemplars and background patches. Furthermore, similar to [58], the authors also proposed a self-attention mechanism to represent self-similarity among features.
@@
Lin et al. [62] focused on the robustness of exemplar feature representations, such as in VCN [59] and RCAC [60], and proposed the Scale-Prior Deformable Convolution Network (SPDCN). The authors argued that most previous works did not consider exemplar scale information during feature extraction or did not do it intuitively. Therefore, they proposed taking advantage of size information provided by the exemplars, discarding features inconsistent with given scales, and consequently extracting more meaningful features. Specifically, they considered a scale-prior backbone consisting of the first ten convolutional layers of an ImageNet pre-trained VGG-19 network [96], where some specific layers have been converted to scale-prior deformable convolutions [98, 99] to exploit the critical scale information from the exemplars. Afterward, these semantic features are split into two branches for segmentation and density estimation, respectively. In the former branch, such as in FamNet and other previous works, ROI pooling [88] is exploited to extract feature maps concerning the exemplars; next, these features are averaged into a representation vector, and a whole similarity map is calculated by computing the cosine similarity between this vector and the image feature map. In the density estimation branch, class-agnostic features are calculated using element-wise multiplication like in [37, 59, 58, 60] between the similarity map and the image feature map. Then, they are fed into a decoder that follows the design of the Pixel Shuffle Crowd Counter (PSCC) [100], which employs pixel-shuffling [101] for the upsampling operations, and that is responsible for predicting the final density map. Finally, another novel contribution of this paper is the usage of the generalized loss [102] to optimize the density predictor. Unlike the MSE loss, utilized in most previous works, generalized loss measures the distance between the predicted density map directly against the dot map, and not the ground truth density map generated convolving Gaussian kernels. In this way, the cost function is adjusted accordingly to the given exemplars, and it is not fixed to the object-specific sizes.
Figure 4: Train and inference high-level overview of reference-based CAC approaches relying on density regression. Following the notation of this survey, CAC models are fed with query images $I_{i}$ and $K$ exemplars expressed as bounding box coordinates $B^{E}=\{b_{i}\}_{i=1:k}\in\mathbb{R}^{4}$. Usually, $K=3$. Exemplars belong to several object classes, which are different at training and test time. The model is in charge of (i) computing feature representations from these inputs that should be agnostic from specific object classes, and (ii) predicting density maps $D^{map}$ from these feature representations. Density-based methods are the standard approach for object counting in crowded scenes. Training typically involves minimizing a per-pixel loss between the predicted and ground-truth density maps, with the final count obtained by summing the pixel values of the predicted density maps.
@@
The authors in [63] argued that bounding boxes are not the ideal representations for most exemplars as they often capture undesirable background portions that harm the matching process. Therefore, they proposed exploiting segmentation masks as an alternative, introducing the Mask Augmented Counting network (MACnet). In more detail, the proposed architecture consists of three modules. First, an ImageNet [103] pre-trained ResNet-50 [87] is used as a feature extractor to embed both the query and the exemplars into a feature space. Specifically, its input consists of the query image $I$, the exemplars bounding box coordinates $B^{E}$, and their associated masked counterparts $B_{m}^{E}$. However, unlike previous works, the exemplars are not retrieved with bounding boxes but through their extreme points. The reason claimed by the authors is twofold: (i) annotation with extreme points is approximately four times faster than using bounding boxes [104], and (ii) extreme points are exploited to compute the segmentation masks with a pre-trained Deep Extreme Cut network (DEXTR) [105]. It is also worth noting that while previous literature got the exemplars by cropping them in the feature space, here they are retrieved in the pixel space since elements in the extracted feature tensor already capture background information due to the large receptive field of the feature extractors. The second module consists of a residual layer fed with the feature representation of the exemplars and their masked counterparts. Specifically, it mitigates errors during the segmentation mask creation, learning from the residual information instead of immediately discarding the background information. Finally, the third module is responsible for predicting the density map, supervised by the MSE loss. As in previous papers [37, 58], it is fed with a similarity map obtained by convolving the feature map of the query image with the feature map resulting from the residual module as its filter, i.e., by performing the inner product between these two feature maps. However, here, the final count is derived not directly from integrating the density map but by counting the peaks utilizing a peak-finding algorithm. A final observation concerns the usage of a test time adaptation loss like in [37, 59] aiming to encourage the network in distinguishing background from foreground in the provided exemplars.
@@
You et al. introduced SAFECount in [64], which uses Similarity-Aware Feature Enhancement blocks. These blocks compose of two main modules; (i) a Similarity Comparison Module (SCM), and (ii) a Feature Enhancement Module (FEM). In SCM, query and reference features extracted by an ImageNet pre-trained ResNet-18 [87] are compared in a learnable projection space, and a score map is obtained for each reference image. The similarity map is obtained from the score map by normalizing the exemplar dimension and spatial dimension. The FEM uses the similarity map to weight the reference features, highlighting regions in the query image that are similar to the reference images. These weighted features are then fused into the query features, creating an enhanced representation that emphasizes the areas containing objects of interest.
@@
Djukic et al. focused on the strength of exemplar feature representations, arguing that, ideally, the prototypes should generalize over the appearance of the selected object category in the image. They introduced the Low-shot Object Counting network with iterative prototype Adaptation (LOCA) [65]. The core of their solution is the object prototype extraction module, which separately analyzes the shape and appearance characteristics of exemplars and iteratively refines them into object prototypes. Specifically, appearance information is injected by initializing the prototypes with image features computed by RoI pooling [88] exemplar bounding boxes over a query image feature map obtained with a pre-trained ResNet50 [87]. However, the pooling operation makes the appearance queries shape-agnostic, since it maps features from different spatial shapes into rectangular queries of the same size. Thus, the authors injected also shape information, by initializing the prototypes with exemplar width and height features. Finally, the shape and appearance queries are converted into object prototypes by an iterative adaptation module using a recursive sequence of cross-attention blocks.
@@
Soliven et al. introduced the Contrast Counting Network (ConCoNet) in [66]. They proposed to use not only the (positive) exemplars as in previous works but also the negative exemplars, i.e., object instances defining what not to count. Their contribution can be incorporated into other class-agnostic counting models, and, by way of examples, they reported the results obtained using two of the milestones for the reference-based class-agnostic counting task, e.g., FamNet [37] and BMNet+ [61]. Specifically, the class-agnostic similarity maps are computed not only for the query image and positive exemplar pairs but also for the negative ones; the final similarity map inputted to the density regressor is obtained by simply subtracting the negative similarity map from the positive one. Since CAC datasets do not contain negative exemplars, the authors derived them under the assumption that users provide negative examples during inference, thus performing a sort of adaptation at test time; instead, to enable the training, they selected bounding boxes randomly from the query image regions classified as background by both the ground-truth and the predicted density maps.
@@
The authors in [70] focused on a novel correlation mechanism between exemplar-query features, overcoming the limitations of kernel-wise convolution characterizing previous works such as FamNet. Specifically, they proposed CountVers, which integrates both channel-wise and spatial-wise correlation mechanisms for simultaneously counting small and large objects. The innovative perspective of this work relies on two key insights. On the one hand, small objects in query images tend to have features concentrated in specific spatial regions, highlighting the importance of the channel dimension and thus suggesting channel-wise correlation mechanisms for enhancing the counting accuracy of small objects. On the other hand, large objects often span significant spatial areas, resulting in complex spatial semantic features, and thus utilizing spatial-wise correlation mechanisms effectively is a crucial factor for ensuring accurate counting of large objects.
@@
Recently, Zhicheng et al. proposed CACViT [71], a framework for CAC that leverages a single pre-trained standard Vision Transformer (ViT). This work highlights the practicality of using a plain ViT for CAC by introducing a novel "extract-and-match" paradigm, which eliminates the need for multiple feature extractors or post-processing for feature matching. Specifically, the authors demonstrated that the attention mechanism of ViT can not only extract features from both the query and exemplar images but also perform feature matching. By concatenating the query and exemplar tokens and inputting them into a basic ViT model, the self-attention process within the ViT is viewed as comprising two distinct groups: one involving self-attention and the other cross-attention. The self-attention component is responsible for extracting features from both the query and exemplar images, while the cross-attention facilitates the matching between them.
@@
Pelhan et al. recently introduced a detect-and-verify paradigm (DAVE) [72] that combines the strengths of density-based methods – best suited for crowded scenarios – with those of detection-based techniques, which provide precise object localization. This approach follows a two-stage pipeline. In the first detection stage, DAVE generates a high-recall set of candidate bounding boxes by first estimating the object centers through a location density map and then predicting bounding box parameters using a regression head. The verification stage refines these detections by removing false positives. Each candidate is analyzed based on its appearance features, which are compared to features of the provided exemplars using clustering methods. Detections that do not align with exemplar clusters are classified as outliers and removed. The refined bounding boxes are then used to update the density map, which further improves count estimation accuracy.
@@
A token-based framework for unified open-world text-guided counting (UPC) is introduced in [74]. Prompts, which may take the form of boxes, points, or text, are first transformed into a unified representation called a prompt mask. For box and point prompts, this mask is derived directly from the labeled regions or pixels within the image. As in previous methods, image features are extracted using a convolutional neural network (CNN) encoder, and a cross-attention mechanism is applied to determine the similarity between the prompt token and the image features. This similarity guides the creation of density features, which are subsequently decoded by a CNN to produce the final density map. However, in this work, to improve the model’s robustness, the predicted density map is iteratively refined by treating it as a new prompt mask in successive iterations. This iterative process establishes a fixed point, allowing the model to converge to consistent predictions. A custom-designed fixed-point loss function, grounded in implicit differentiation, further stabilizes training while optimizing parameters effectively. An additional distinctive aspect of this approach is the implementation of a contrastive training strategy. This strategy pairs positive samples, which contain objects matching the prompt, with negative samples, which do not. By compelling the model to generate accurate density maps for positive samples and zero-density maps for negative samples, this method significantly reduces errors associated with noisy or ambiguous prompts.
@@
[75] As in TFPOC [73], the authors in [75] exploited a detection-based approach relying on the popular Segment Anything Model [107] for instance segmentation. Specifically, they introduced a framework termed Point, Segment and Count (PseCo). Different from TFPOC, PseCo does not use uniform grid points to prompt the Segment Anything Model to segment all objects. Indeed, the authors argued that using grid points may fall short in capturing all objects, particularly in densely crowded scenes, leading to the omission of many small objects and making this approach unsuitable for object counting. Although increasing the density of grid points could help address this limitation, it would result in substantial computational overhead, as a significant number of points would still target background regions. This makes such an approach impractical for real-world use. Thus, in [75] a class-agnostic object localization that estimates a heatmap of all objects was proposed. From this heatmap, object point coordinates are inferred and are used to feed the Segment Anything Model. Following this step, PseCo introduced a second stage that leveraged another foundation model alongside the Segment Anything Model, namely the popular Contrastive Language–Image Pretraining (CLIP) model [108]. Specifically, CLIP is used to classify the regions found in the first stage with respect to the given exemplars. To encourage a discriminative classification among mask proposals, a novel hierarchical knowledge distillation between mask proposals was also proposed.

### v4

@@ III-B Reference-based Class-agnostic Counting
The seminal work on reference-based object counting is by Ranjan et al. [32], who first formulated counting as a few-shot regression task. Since the task was novel, the authors also introduced the first dataset tailored for it, a collection of images containing objects belonging to 147 categories named FSC-147; labels comprise dots for the training stage, and exemplar bounding boxes that are used both in the supervised learning and at inference time. See also Sec. IV for more details. Finally, the authors proposed the Few-Shot Adaptation and Matching Network (FamNet), a novel architecture designed to address this task. Specifically, it includes two main modules: (i) a multi-scale feature extraction module and (ii) a density prediction module. The first consists of the first four blocks picked up from an ImageNet pre-trained ResNet-50 backbone [89], frozen during training. The convolutional feature maps produced by the third and fourth blocks provide the representation of the images $I_{i}$; furthermore, they are also exploited to obtain the multi-scale features of the exemplar images obtained from $B^{E}$ by performing ROI pooling [90] at different scales. Rather, the density prediction module consists of some stacked convolution blocks and upsampling layers; the last one is a $1\times 1$ convolutional layer in charge of predicting the final $2D$ density map $D^{map}$ having the same size as the input image. The input of this second module does not correspond to the features produced by the feature extractor, as usual in the class-specific counting task. Instead, the authors computed the convolution operation between the scaled exemplar features and the image features, obtaining multiple correlation maps, one for each scale, that are then concatenated and fed to the density predictor. The reason behind this is making the density regressor agnostic to the visual categories. During the learning stage, FamNet is trained by minimizing the mean squared error (MSE) between the predicted and the ground truth density map. Finally, this work also proposed a sort of fine-tuning that exploits test data to improve performance further. Specifically, it introduced an adaptation loss that exploits the locations of the exemplars included in the test set to be used only at inference time. It comprises two loss components: (i) the Min-Count loss, ensuring that the density sum within each exemplar box is $\geq$ 1, and (ii) the perturbation loss, encouraging Gaussian-shaped density around exemplars—similar to correlation filters used in tracking [91, 92, 93].
@@
Roughly simultaneously, Yang et al. introduced a similar solution [60] named Class-agnostic Few-shot Object Counting Network (CFOCNet). The architecture follows the scheme of FamNet, with an encoder and decoder. The encoder includes two ResNet-50 streams for feature extraction: one for the query image and one for the exemplars (here referred to as reference images). The features concerning the exemplars are matched by computing the convolution operation with the output of a self-attention mechanism [94] applied to the features relating to the query image. These resulting matching maps are finally fused by exploiting a scale-aware fusing mechanism, and the output is fed into the decoder to compute the density map. The decoder is similar to the one exploited in FamNet—some stacked convolution blocks and upsampling layers—and it is supervised by standard MSE loss as well as by SSIM loss [95] to catch local pattern consistency [96].
@@
Some of the authors of [32] extended FamNet in [61], addressing its limitations due to the small dataset used for supervised training. They introduced the Vicinal Counting Network (VCN), which retains the core of FamNet with minor changes and incorporates a generator to augment and synthesize images from the vicinity of training samples, expanding the dataset. Specifically, the generator is fed with an input image $I\in\mathbb{R}^{H\times W\times 3}$ and a noise vector $z\in\mathbb{R}^{m}$, and it produces an augmented version $I_{z}\in\mathbb{R}^{H\times W\times 3}$ of $I$. This augmented image serves as the input to the feature extractor and the subsequent regressor module, together with the original input $I$. The entire system is then trained jointly in a cooperative fashion. However, unlike the min-max game typical of GANs [97], VCN relies on a standard minimization problem, combining MSE loss for density regression with a reconstruction loss to enforce similarity between generated and input samples, and a diversity loss to prevent identity mapping from $I$ to $I_{z}$. Finally, the authors used the same test time adaptation loss of FamNet [32].
@@
Another work inspired by FamNet is [62]. First, the authors explored the main limitations of FamNet by analyzing failure cases, identifying a strong dependence on exemplar diversity—exemplar boxes are few and subjectively annotated, and they may not cover all object instances or ensure intra-class diversity. To address this, they proposed the Robust Class-Agnostic Counter (RCAC), which uses a two-stream architecture: one branch processes RGB images $I_{i}$ with a FamNet-like feature extractor, while the other uses a VGG-based extractor on edge images derived from $I_{i}$ throughout an edge detector. This second feature extractor is randomly initialized and, different from the RGB feature extractor, optimized during training. The reason claimed by the authors behind using edge images is that edges are a sort of class-agnostic knowledge, i.e., shape is a more reliable cue across instances than color. Then, ROI pooling is applied to both branches to crop exemplar features, which are then enhanced via a feature augmentation module. After that, feature correlation is obtained by computing the convolution operation between the latter and the full feature maps, like in previous works, and the density map is predicted using the same module as in FamNet. In detail, the feature augmentation module aims to generate more exemplars of different colors, shapes, and scales in the feature space by combining the provided exemplars, sharing the goal of VCN [61] to expand the training set.
Figure 4: Train and inference high-level overview of reference-based CAC approaches relying on density regression. Following the notation of this survey, CAC models are fed with query images $I_{i}$ and $K$ exemplars expressed as bounding box coordinates $B^{E}=\{b_{i}\}_{i=1:k}\in\mathbb{R}^{4}$. Usually, $K=3$. Exemplars belong to several object classes, which are different at training and test time. The model is in charge of (i) computing feature representations from these inputs that should be agnostic to specific object classes, and (ii) predicting density maps $D^{map}$ from these feature representations. Density-based methods are the standard approach for object counting in crowded scenes. Training typically involves minimizing a per-pixel loss between the predicted and ground-truth density maps, with the final count obtained by summing the pixel values of the predicted density maps.
@@
Another notable work that represents a milestone is [63]. Here, Shi et al. argued that two factors played crucial roles in previous works: (i) feature representation of query images and exemplars, and (ii) similarity metric for matching exemplar and image features. Previous methods either exploited a fixed [32, 61] or a learnable [60] feature extractor but employed a similarity metric relying on the fixed inner product [32, 61, 60, 62, 65], which yielded insufficient matching outcomes. Hence, the authors proposed the Bilinear Matching Network (BMNet+) to improve the similarity modeling. Specifically, the authors designed a learnable bilinear similarity loss to supervise the similarity-matching results, thus going beyond the fixed inner product. The authors took inspiration from metric learning, which aims to embed data into a space where similar samples are pulled close while dissimilar ones are pushed away [98]. In more detail, their similarity loss pulls close the features between the exemplar and target instances while pushing away the features between exemplars and background patches.
@@
Lin et al. [64] focused on the robustness of exemplar feature representations, such as in VCN [61] and RCAC [62], and proposed the Scale-Prior Deformable Convolution Network (SPDCN). The authors argued that most prior works often overlooked or poorly handled the exemplar scale during feature extraction. To address this, they leveraged size information from exemplars by exploiting scale-prior deformable convolutions [99, 100] to discard features inconsistent with scale, enabling more meaningful feature extraction. Afterward, these semantic features are split into two branches for segmentation and density estimation, respectively. In the former branch, such as in FamNet and other previous works, ROI pooling [90] is exploited to extract feature maps concerning the exemplars; next, these features are averaged into a representation vector, and a whole similarity map is calculated by computing the cosine similarity between this vector and the image feature map. In the density estimation branch, class-agnostic features are calculated using element-wise multiplication like in [32, 61, 60, 62] between the similarity map and the image feature map. Then, they are fed into a decoder that follows the design of the Pixel Shuffle Crowd Counter (PSCC) [101], which employs pixel-shuffling [102] for the upsampling operations, and that is responsible for predicting the final density map.
@@
The authors in [65] argued that bounding boxes are not the ideal representations for most exemplars as they often capture undesirable background portions that harm the matching process. Therefore, they proposed exploiting segmentation masks as an alternative, introducing the Mask Augmented Counting network (MACnet). In more detail, the proposed architecture consists of three modules. First, an ImageNet [103] pre-trained ResNet-50 [89] is used as a feature extractor to embed both the query and the exemplars into a feature space. Specifically, its input consists of the query image $I$, the exemplars’ bounding box coordinates $B^{E}$, and their associated masked counterparts $B_{m}^{E}$ computed with the pre-trained Deep Extreme Cut network (DEXTR) [104]. The second module consists of a residual layer fed with the feature representation of the exemplars and their masked counterparts. Specifically, it mitigates errors during segmentation mask creation by learning from residual information instead of immediately discarding background information. Finally, the third module is responsible for predicting the density map, supervised by the MSE loss. As in previous papers [32, 60], it is fed with a similarity map obtained by convolving the feature map of the query image with the feature map resulting from the residual module as its filter, i.e., by performing the inner product between these two feature maps. However, here, the final count is derived not directly from integrating the density map but by counting the peaks utilizing a peak-finding algorithm. A final observation concerns the usage of a test time adaptation loss like in [32, 61], aiming to encourage the network in distinguishing background from foreground in the provided exemplars.
@@
You et al. introduced SAFECount in [66], which uses Similarity-Aware Feature Enhancement blocks. These blocks are composed of two main modules: (i) a Similarity Comparison Module (SCM), and (ii) a Feature Enhancement Module (FEM). In SCM, query and reference features extracted by an ImageNet pre-trained ResNet-18 [89] are compared in a learnable projection space, and a score map is obtained for each reference image. The similarity map is obtained from the score map by normalizing the exemplar dimension and spatial dimension. The FEM uses the similarity map to weight the reference features, highlighting regions in the query image that are similar to the reference images. These weighted features are then fused into the query features, creating an enhanced representation that emphasizes the areas containing objects of interest.
@@
Djukic et al. focused on the strength of exemplar feature representations, arguing that, ideally, the prototypes should generalize over the appearance of the selected object category in the image. They introduced the Low-shot Object Counting network with iterative prototype Adaptation (LOCA) [67]. The core of the method is an object prototype extraction module that separately encodes shape and appearance from exemplars and refines them iteratively. Appearance is injected by initializing prototypes with RoI-pooled features from exemplar boxes over a ResNet50-based query feature map. However, this pooling makes the queries shape-agnostic, so shape information is also injected using exemplar width and height. Then, the final prototypes are obtained via an iterative adaptation module with cross-attention blocks.
@@
Soliven et al. introduced the Contrast Counting Network (ConCoNet) in [68]. They proposed to use not only the (positive) exemplars as in previous works but also the negative exemplars, i.e., object instances defining what not to count. Their contribution can be incorporated into other class-agnostic counting models, and, by way of examples, they reported the results obtained using two of the milestones for the reference-based class-agnostic counting task, e.g., FamNet [32] and BMNet+ [63]. Specifically, the class-agnostic similarity maps are computed not only for the query image and positive exemplar pairs but also for the negative ones; the final similarity map input to the density regressor is obtained by simply subtracting the negative similarity map from the positive one. Since CAC datasets do not contain negative exemplars, the authors derived them under the assumption that users provide negative examples during inference, thus performing a sort of adaptation at test time; instead, to enable the training, they selected bounding boxes randomly from the query image regions classified as background by both the ground-truth and the predicted density maps.
@@
The authors in [72] focused on a novel correlation mechanism between exemplar-query features, overcoming the limitations of kernel-wise convolution characterizing previous works such as FamNet. Specifically, they proposed CountVers, which integrates both channel-wise and spatial-wise correlation to handle small and large object counting. The method builds on two insights: small objects have localized features, making channel-wise correlation beneficial, while large objects span broader regions with complex spatial semantics, requiring spatial-wise correlation for accurate counting.
@@
Recently, Zhicheng et al. [73] introduced a framework based on a pre-trained Vision Transformer (ViT) and an "extract-and-match" paradigm that removes the need for multiple feature extractors or post-processing. The attention mechanism in ViT is shown to handle both feature extraction and matching by processing concatenated query and exemplar tokens. Within the ViT, self-attention extracts features from both inputs, while cross-attention enables matching between them.
@@
Pelhan et al. recently introduced a detect-and-verify paradigm (DAVE) [74] that combines the strengths of density-based methods—best suited for crowded scenarios—with those of detection-based techniques, which provide precise object localization. This approach follows a two-stage pipeline. In the first detection stage, DAVE generates a high-recall set of candidate bounding boxes by first estimating the object centers through a location density map and then predicting bounding box parameters using a regression head. The verification stage refines these detections by removing false positives. Each candidate is analyzed based on its appearance features, which are compared to features of the provided exemplars using clustering methods. Detections that do not align with exemplar clusters are classified as outliers and removed. The refined bounding boxes are then used to update the density map, which further improves count estimation accuracy.
@@
A token-based framework for unified open-world text-guided counting (UPC) is introduced in [76]. Prompts, which may take the form of boxes, points, or text, are first transformed into a unified representation called a prompt mask. For box and point prompts, this mask is derived directly from the labeled regions or pixels within the image. As in previous methods, image features are extracted using a CNN-based encoder, and a cross-attention mechanism is applied to determine the similarity between the prompt token and the image features. This similarity guides the creation of density features, which are subsequently decoded by a CNN to produce the final density map. However, in this work, to improve the robustness of the model, the predicted density map is iteratively refined by reusing it as a new prompt mask, leading to convergence toward consistent outputs. A fixed-point loss function, based on implicit differentiation, stabilizes training and improves parameter optimization. Additionally, a contrastive training strategy is used: positive samples contain objects matching the prompt, while negative ones do not. This forces the model to produce accurate density maps for positives and zero-density maps for negatives, reducing errors from noisy or ambiguous prompts.
@@
As in TFPOC [75], the authors in [77] exploited a detection-based approach relying on the popular SAM [106] for instance segmentation. Specifically, they introduced a framework termed Point, Segment and Count (PseCo). Different from TFPOC, PseCo does not use uniform grid points to prompt SAM to segment all objects. Indeed, the authors argued that using grid points may fall short in capturing all objects, particularly in densely crowded scenes, leading to the omission of many small objects. Although increasing the density of grid points could help address this limitation, it would result in substantial computational overhead. Thus, a class-agnostic object localization method was proposed in [77], estimating a heatmap from which object coordinates are inferred and used to guide SAM. then introduced a second stage that integrates another foundation model, the Contrastive Language–Image Pretraining (CLIP) model [107], to classify regions identified in the first stage based on given exemplars.
CountDiff
Hui et al. [78] proposed CountDiff, a framework designed to boost CAC by leveraging knowledge from a pre-trained diffusion model. To encode exemplar-specific information, CountDiff learns an object-specific embedding by mapping image features into the text embedding space, enabling the diffusion model to interpret the visual characteristics of the target object. To enhance generalization in CAC, CountDiff complements this with an object-agnostic embedding, capturing transferable knowledge from the diffusion model’s cross-attention layers. By combining both embeddings, the framework effectively exploits cross-attention maps to identify regions relevant for CAC. Finally, a lightweight test-time adaptation step fine-tunes the object-specific embedding using the provided exemplars, further boosting accuracy on novel objects.